### v1

# Bayesian Multivariable Bidirectional Mendelian Randomization

Abstract
Mendelian randomization (MR) is a pivotal tool in genetic epidemiology, leveraging genetic variants as instrumental variables to infer causal relationships between modifiable exposures and health outcomes. Traditional MR methods, while powerful, often rest on stringent assumptions such as the absence of feedback loops, which are frequently violated in complex biological systems. In addition, many popular MR approaches focus on only two variables (i.e., one exposure and one outcome) whereas our motivating applications have many variables. In this article, we introduce a novel Bayesian framework for multivariable MR that concurrently addresses unmeasured confounding and feedback loops. Central to our approach is a sparse conditional cyclic graphical model with a sparse error variance-covariance matrix. Two structural priors are employed to enable the modeling and inference of causal relationships as well as latent confounding structures. Our method is designed to operate effectively with summary-level data, facilitating its application in contexts where individual-level data are inaccessible, e.g., due to privacy concerns. It can also account for horizontal pleiotropy. Through extensive simulations and applications to the GTEx and OneK1K data, we demonstrate the superior performance of our approach in recovering biologically plausible causal relationships in the presence of possible feedback loops and unmeasured confounding. The R package that implements the proposed method is available at MR.RGM.
Keywords: Causal Inference, Instrumental Variables, Unmeasured Confounding, Summary-Level Data, Gene Regulatory Networks, Horizontal Pleiotropy.
1 Introduction
Mendelian randomization (MR) is a causal inference framework using genetic variants as instrumental variables and has revolutionized the fields of genetics and epidemiology. The main principle of MR is rooted in Mendel’s laws of inheritance, which ensure the random allocation of alleles, thereby mitigating confounding and reverse causation that often plague observational studies davey2003 . This natural randomization resembles the design of randomized controlled trials, offering a powerful alternative when such trials are infeasible.
To date, a wide range of MR methods have been developed including classic methods such as inverse-variance weighting (IVW burgess2013mendelian ) for summary-level data and the two-stage least squares basmann1957generalized ; theil1953repeated for individual-level data as well as more modern approaches such as MR-Egger regression bowden2015mendelian , the weighted median estimator bowden2016 , MR-PRESSO verbanck2018 , the weighted mode-based estimator hartwig2017 , and the generalized summary-data-based Mendelian randomization zhu2018 .
There is also a range of software packages for MR. The MendelianRandomization package R-MendelianRandomization provides R implementations of the IVW, MR-Egger, simple and weighted median, and intercept-based estimators. The TwoSampleMR package R-TwoSampleMR ; Hemani2018 enables systematic two-sample MR analyses using summary-level data, offering a suite of MR methods along with data harmonization tools. The OneSampleMR package R-OneSampleMR is designed for analyses using individual-level data, supporting two-stage predictor substitution and two-stage residual inclusion approaches, which are appropriate when genetic instruments and exposures are measured within the same dataset.
Despite the large literature of MR methods and software packages, the focus has been predominantly on the “one exposure and one outcome” setting, under which only the total causal effect can be inferred. Falling short in addressing the complexities of multiple exposures and outcomes, these MR methods cannot differentiate between direct and indirect/mediation effects. Recognizing the complex nature of biological systems, multivariable MR (MVMR) was developed to estimate the direct effects of multiple exposures on an outcome simultaneously sanderson2019 . MVMR accounts for the correlation between exposures, providing a more nuanced understanding of causal pathways. The mrbayes package R-MrBayes provides Bayesian implementations of IVW and MR-Egger for two-sample MR, including multivariable extensions. GRIVET is a recent approach that infers causal relationships among a set of variables in the presence of unmeasured confounding by leveraging directed acyclic graph (DAG) models chen2023discovery . MrDAG zuber2025bayesian is a Bayesian DAG model utilizing genetic variants as instruments.
While these MVMR approaches have enabled richer causal inference by discerning direct and indirect causal effects, their acyclic assumption excludes the possibility of directed cycles or feedback loops, which are prevalent in many biological systems such as gene regulatory networks, metabolic networks, and phenotypic disease networks. To capture potential feedback loops, Spirtes (1995) spirtes1995directed extends DAG models to directed cyclic graph (DCG) models, which have been applied to genomics cai2013inference ; ni2018reciprocal ; ni2018heterogeneous , brain imaging roy2023directed , and electronic health records jin2025directed . However, they all assume there is no unmeasured confounding and hence would draw biased inference when the assumption is violated.
In this paper, we propose a new Bayesian MVMR approach based on non-recursive structural equation models (SEMs) with sparse correlated errors, termed Mendelian randomization with reciprocal graphical model (MR.RGM), which simultaneously addresses unmeasured confounding and feedback loops using genetic variants as instrumental variables. Besides addressing the challenges of unmeasured confounding and feedback loops, our approach has four additional features. First, it does not require individual-level data; only summary-level data would suffice. In many MR studies, the genetic variants are not publicly available, and only summary statistics are published. Our approach is applicable to those common settings. Second, our approach also infers the exact positions of confounding via a graphical spike-and-slab prior, i.e., the identification of which sets of variables are confounded. Third, the Bayesian nature of our approach allows for natural uncertainty quantification and enhanced interpretability by providing credible intervals for causal effects and posterior inclusion probabilities. Finally, we also extend MR.RGM to account for horizontal pleiotropy and weak instruments.
To assess the performance of our approach, we conduct extensive simulation studies with realistic network structures such as scale-free networks and small-world networks, comparing its performance against existing MR methods that are implemented in the popular MR software packages, MendelianRandomization, mrbayes, and OneSampleMR. Furthermore, we apply the poposed method to two real-world datasets, the GTEx V7 skeletal muscle data gtex2017genetic and the OneK1K data onek1k2022 , successfully recovering biologically plausible causal relationships supported by existing literature.
2 Method
2.1 MR, Bidirectional MR, and MVMR
MR.
MR is an instrumental variable approach to infer causal relationships between exposures and outcomes using genetic variants as instruments. The validity of MR rests on three key assumptions: (i) the genetic variants must be associated with the exposure of interest (relevance); (ii) they must be independent of any confounders of the exposures and outcomes (independence); and (iii) they should influence the outcome only through their effect on the exposure, without any alternative pathways (exclusion restriction). Traditionally, MR only considers one exposure $Y_{1}$ and one outcome $Y_{2}$ with the direction of causality fixed to be $Y_{1}\to Y_{2}$. Using $X_{1}$ as an instrument for the exposure $Y_{1}$, MR considers the following generative model,
$Y_{1}=b_{1}X_{1}+c_{1}W+E_{1}$ (1)
$Y_{2}=a_{21}Y_{1}+c_{2}W+E_{2},$ (2)
where $W$ is an unmeasured confounder of $Y_{1}$ and $Y_{2}$, and $E_{1},E_{2}$ are independent exogenous errors. The confounder induces non-causal association between $Y_{1}$ and $Y_{2}$, which in turn leads to a biased estimate of $a_{21}$, the main quantity of interest, if we simply regress $Y_{2}$ on $Y_{1}$. Equations (1)-(2) as a generative model imply the exclusion restriction as well as the relevance assumption (as long as $b_{1}\neq 0$). We additionally assume $X_{1}\perp W$ (i.e., the independence assumption).
Two-stage least squares (2SLS) and inverse-variance weighting (IVW) are the two most commonly used methods to estimate the causal effect $a_{21}$ of exposure $Y_{1}$ on outcome $Y_{2}$. In 2SLS, we first regress $Y_{1}$ on $X_{1}$ and get a fitted $\widehat{Y}_{1}$ and then regress $Y_{2}$ on $\widehat{Y}_{1}$. The coefficient of $\widehat{Y}_{1}$ in the latter regression is the desired causal effect $a_{21}$. Such an approach can be justified by plugging (1) into (2),
$Y_{2}$ $=a_{21}b_{1}X_{1}+a_{21}c_{1}U+a_{21}E_{1}+c_{2}W+E_{2}$ (3)
$=a_{21}\widehat{Y}_{1}+a_{21}c_{1}U+a_{21}E_{1}+c_{2}W+E_{2},$ (4)
where the second equality uses the fact that $\widehat{Y}_{1}=b_{1}X_{1}$ at the population level because $X_{1}$ is independent of $W$ and $E_{1}$. This independence (plus the independence of $X_{1}$ and $E_{2}$) and (4) together imply that $a_{21}$ can be recovered by regressing $Y_{2}$ on $\widehat{Y}_{1}$.
IVW relies on the fact that
$a_{21}=\frac{a_{21}b_{1}}{b_{1}}:=\frac{r_{21}}{r_{11}}$ (5)
where the numerator $r_{21}$ is the effect of $X_{1}$ on $Y_{2}$ in a simple linear regression model due to (3) and the mutual independence of $X_{1},W,E_{1},E_{2}$, and the denominator $r_{11}$ is the effect of $X_{1}$ on $Y_{1}$ in a simple linear regression model due to (1) and the same mutual independence assertion. IVW then takes a weighted average of those ratio estimators when there are multiple instruments with weights equal to the inverse variances of the estimators.
Bidirectional MR.
When the direction of causality between $Y_{1}$ and $Y_{2}$ is unknown, separately applying MR in both directions is often adopted, provided that an instrument $X_{2}$ is also available for $Y_{2}$, which is known as the bidirectional MR. When the causal effects in both directions are significant, both may be reported simultaneously, indicating bidirectional/reciprocal causal effects. The validity of using 2SLS and IVW in such a scenario can be justified by considering a non-recursive SEM as a generative model,
$\left\{\begin{array}[]{cc}Y_{1}=a_{12}Y_{2}+b_{1}X_{1}+c_{1}W+E_{1}\\ Y_{2}=a_{21}Y_{1}+b_{2}X_{2}+c_{2}W+E_{2}.\end{array}\right$. (8)
Crucially, unlike Equations (1) and (2), the two equations in (8) are coupled. The right-hand sides of these equations are not the conditional expectations of the left-hand sides unless $a_{12}$ or $a_{21}$ is zero. Under generative model (8), the marginal distributions of $Y_{1}$ and $Y_{2}$ are:
$Y_{1}=\frac{1}{1-a_{12}a_{21}}\left\{b_{1}X_{1}+a_{12}b_{2}X_{2}+(c_{1}+a_{12}c_{2})W+E_{1}+a_{12}E_{2}\right\},$
$Y_{2}=\frac{1}{1-a_{12}a_{21}}\left\{b_{2}X_{2}+b_{1}a_{21}X_{1}+(c_{2}+a_{21}c_{1})W+E_{2}+a_{21}E_{1}\right\}.$ (9)
Because $X_{1},X_{2},W,E_{1},E_{2}$ are mutually independent, we have
$\left\{\begin{array}[]{c}\frac{b_{1}}{1-a_{12}a_{21}}=r_{11}\\ \frac{b_{1}a_{21}}{1-a_{12}a_{21}}=r_{21}\\ \frac{a_{12}b_{2}}{1-a_{12}a_{21}}=r_{12}\\ \frac{b_{2}}{1-a_{12}a_{21}}=r_{22}\end{array}\right$.
where $r_{jk}$ is effect of $X_{k}$ on $Y_{j}$ in a simple linear regression for $j,k\in\{1,2\}$. Hence, the causal effect of $Y_{1}$ on $Y_{2}$ is
$a_{21}=\frac{r_{21}}{r_{11}},$
which coincides with (5) because the extra factor $\frac{1}{1-a_{12}a_{21}}$ due to the coupling in SEM is common to both $r_{21}$ and $r_{11}$ and thus cancels out in the ratio. Similarly, $a_{12}=\frac{r_{12}}{r_{22}}$. Consequently, both 2SLS and IVW are still valid estimation procedures for bidirectional MR.
MVMR.
However, most real-world systems have more than two variables. Consider, for example, a trivariate generative SEM,
$\left\{\begin{array}[]{cc}Y_{1}=a_{12}Y_{2}+a_{13}Y_{3}+b_{1}X_{1}+c_{1}W+E_{1}\\ Y_{2}=a_{21}Y_{1}+a_{23}Y_{3}+b_{2}X_{2}+c_{2}W+E_{2}\\ Y_{3}=a_{31}Y_{1}+a_{32}Y_{2}+b_{3}X_{3}+c_{3}W+E_{3}.\end{array}\right$. (13)
where $a_{jk}$ is the direct causal effect of $Y_{k}$ on $Y_{j}$, often represented graphically by $Y_{k}\to Y_{j}$ or simply $k\to j$ in a causal graph. Unlike the bivariate case, applying MR or bidirectional MR to $(Y_{1},Y_{2})$ does not estimate the direct causal effect $a_{12}$. Instead, it targets the total causal effect $t_{12}$ of $Y_{2}$ on $Y_{1}$, which consists of both direct and indirect/mediated effects via $Y_{3}$ and can be found by do-calculus (pearl1995causal, ):
$t_{12}=\frac{a_{12}+a_{13}a_{32}}{|1-a_{13}a_{31}|},$ (14)
where $a_{13}a_{32}$ is the indirect causal effect and $|1-a_{13}a_{32}|$ is the “amplification” of the causal effect due to the reciprocal causal relationship between $Y_{1}$ and $Y_{3}$. Because $t_{12}$ is generally not equal to $a_{12}$, naively applying MR to construct causal graphs could lead to many false discoveries. For example, consider a hypothetical genetic regulatory cascade $Y_{1}\to Y_{2}\to Y_{3}$. Pairwise applications of MR would impute a false edge $Y_{1}\to Y_{3}$ since the total effect of $Y_{1}$ on $Y_{3}$ is non-zero. For MVMR, a principled approach is the use of graphical models, which aim to estimate the direct causal effects directly. However, as reviewed in Section 1, there is a lack of MVMR methods that can accommodate and estimate reciprocal causality under unmeasured confounding.
2.2 Proposed MR.RGM
Let $\mathbf{Y}=(Y_{1},\cdots,Y_{p})^{T}$ denote $p$ traits, and let $\mathbf{X}=(X_{1},\cdots,X_{k})^{T}$ represent $k$ instrumental variables. In our later applications, the traits are gene expressions, and the instruments are (cis-) single-nucleotide polymorphisms (SNPs) that are significantly correlated with the traits. Let $\mathbf{U}=(U_{1},\cdots,U_{l})^{T}$ denote a set of $l$ covariates. We model the data-generative process by an SEM:
$\mathbf{Y}=\mathbf{A}\mathbf{Y}+\mathbf{B}\mathbf{X}+\mathbf{C}\mathbf{U}+\mathbf{D}\mathbf{W}+\mathbf{E},$ (15)
where $\mathbf{A}=(a_{jh})\in\mathbb{R}^{p\times p}$ with $a_{jh}$ being the direct causal effect of trait $h$ on trait $j$, $\mathbf{B}=(b_{jh})\in\mathbb{R}^{p\times k}$ with $b_{jh}$ capturing the effect of instrumental variable $h$ on trait $j$, $\mathbf{C}=(c_{jh})\in\mathbb{R}^{p\times l}$ with $c_{jh}$ representing the effect of covariate $h$ on trait $j$, $\mathbf{W}=(W_{1},\cdots,W_{t})^{T}$ represents $t$ latent confounders, which are assumed to be $\mathbf{W}\sim\text{N}_{t}(0,\mathbf{I}_{t})$, i.e., $t$ independent sources of unmeasured confounding, $\mathbf{D}=(d_{jh})\in\mathbb{R}^{p\times t}$ with $d_{jh}$ being the impact of unobserved confounder $h$ on trait $j$, and $\mathbf{E}\sim\text{N}_{p}(0,\mathbf{\Sigma})$ with diagonal $\mathbf{\Sigma}$ is the independent exogenous errors. We further assume that there are no self-loops, i.e., $\text{diag}(\mathbf{A})=\mathbf{0}$ and that $\mathbf{X}$, $\mathbf{U}$, $\mathbf{W}$, and $\mathbf{E}$ are all mutually independent. Structural zeros are imposed on $\mathbf{B}$ such that $b_{jh}\neq 0$ if and only if $X_{h}$ is the instrument for $Y_{j}$; this will be later relaxed to account for horizontal pleiotropy and weak instruments.
The presence of the latent confounders $\mathbf{W}$ induces correlation among traits. Define:
$\mathbf{E^{*}}:=\mathbf{D}\mathbf{W}+\mathbf{E}\sim\text{N}_{p}(0,\mathbf{D}\mathbf{D}^{T}+\mathbf{\Sigma})=\text{N}_{p}(0,\mathbf{\Sigma^{*}}),$
where $\mathbf{W}$ has been integrated out. The covariance matrix $\mathbf{\Sigma^{*}}$ of this new error term is not diagonal and hence correlated. A non-zero entry $\mathbf{\Sigma}^{*}_{jh}\neq 0$ for $j\neq h$ indicates the presence of a latent confounder affecting both $Y_{j}$ and $Y_{h}$ as it must be exist some $s$ such that neither $d_{js}$ nor $d_{hs}$ is zero. This allows us to infer potential latent confounding structures directly from the covariance matrix of the errors, without having to assume a known number of confounders.
Rewriting (15) with $\mathbf{E^{*}}$, we have,
$(\mathbf{I}_{p}-\mathbf{A})\mathbf{Y}=\mathbf{B}\mathbf{X}+\mathbf{C}\mathbf{U}+\mathbf{E^{*}}$
$\implies$ $\mathbf{Y}=(\mathbf{I}_{p}-\mathbf{A})^{-1}\mathbf{B}\mathbf{X}+(\mathbf{I}_{p}-\mathbf{A})^{-1}\mathbf{C}\mathbf{U}+(\mathbf{I}_{p}-\mathbf{A})^{-1}\mathbf{E^{*}}.$ (16)
This formulation accommodates feedback loops through $\mathbf{A}$ (e.g., if $a_{jh}\neq 0$ and $a_{hj}\neq 0$, then $Y_{j}\rightleftarrows Y_{h}$) and models latent confounding via the error covariance structure. The conditional distribution of $\mathbf{Y}$ given $\mathbf{X}$ and $\mathbf{U}$ can be derived from (2.2):
$\mathbf{Y}\mid\mathbf{X},\mathbf{U}\sim\text{N}_{p}\left\{(\mathbf{I}_{p}-\mathbf{A})^{-1}\mathbf{B}\mathbf{X}+(\mathbf{I}_{p}-\mathbf{A})^{-1}\mathbf{C}\mathbf{U},\,(\mathbf{I}_{p}-\mathbf{A})^{-1}\mathbf{\Sigma^{*}}(\mathbf{I}_{p}-\mathbf{A})^{-T}\right\}.$ (17)
To enable the use of summary-level data—common in MR where individual-level data may be unavailable due to privacy concerns—we represent the conditional distribution in (17) in terms of sufficient statistics, which are empirical second-moment matrices:
$\mathbf{S_{yy}}$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbf{y_{i}}\mathbf{y_{i}}^{T},\quad\mathbf{S_{yx}}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{y_{i}}\mathbf{x_{i}}^{T},\quad\mathbf{S_{yu}}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{y_{i}}\mathbf{u_{i}}^{T},$
$\mathbf{S_{xx}}$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbf{x_{i}}\mathbf{x_{i}}^{T},\quad\mathbf{S_{uu}}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{u_{i}}\mathbf{u_{i}}^{T},\quad\mathbf{S_{xu}}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{x_{i}}\mathbf{u_{i}}^{T}$
These sufficient statistics allow us to make causal inference without requiring access to individual-level data. The distribution (17), based on the sufficient statistics, is given by:
$\begin{split}&p\left(\{\mathbf{y_{i}}\}_{i=1}^{n}|\{\mathbf{x_{i}}\}_{i=1}^{n},\{\mathbf{u_{i}}\}_{i=1}^{n},\mathbf{A,B,C,\Sigma^{*}}\right)\\ =&(2\pi)^{-\frac{np}{2}}.\mathbf{\text{det}(\Sigma^{*})^{-\frac{n}{2}}.|\text{det}(I_{p}-A)|^{n}}.\exp(-\frac{1}{2}n.\mathbf{[\text{tr}(S_{yy}(I_{p}-A)^{T}{\Sigma^{*}}^{-1}(I_{p}-A))}\\ &\quad\quad\quad\quad\quad\quad-2n.\mathbf{\text{tr}(S_{yx}B^{T}{\Sigma^{*}}^{-1}(I_{p}-A))+n.\text{tr}(S_{xx}B^{T}{\Sigma^{*}}^{-1}B)}\\ &\quad\quad\quad\quad\quad\quad-2n.\text{tr}\mathbf{(S_{yu}^{T}C^{T}{\Sigma^{*}}^{-1}(I_{p}-A))-2n.\text{tr}(S_{xu}C^{T}{\Sigma^{*}}^{-1}B)}\\ &\quad\quad\quad\quad\quad\quad+n.\text{tr}\mathbf{(S_{uu}C^{T}{\Sigma^{*}}^{-1}C)]}).\end{split}$
Our goal is to estimate the matrices $\mathbf{A},\mathbf{B},\mathbf{C}$, and $\mathbf{\Sigma^{*}}$, which together capture the causal relationships among traits, instrumental effects, covariate effects, and the confounding effects. To achieve this, we adopt a fully Bayesian framework: we place prior distributions on these parameters and use Markov chain Monte Carlo (MCMC) to sample them from their joint posterior distribution. Using sufficient statistics also improves the scalability with respect to the sample size $n$. The cost of evaluating the likelihood based on raw data is $\mathcal{O}\!\big(n\{p^{2}+p(k+l)\}+p^{3}\big)$ whereas that of sufficient statistics is $\mathcal{O}\!\big(p^{3}+p^{2}(k+r)+p(k+l)^{2}\big)$, which is a big reduction if $n\gg\{p,k,l\}$. In the following sections, we detail the prior specifications and the posterior inference procedure.
2.3 Priors and Posteriors
The priors are chosen to support both parameter estimation and structural learning, with a particular emphasis on inducing sparsity in the causal graph (via $\mathbf{A}$) and in the confounding structure (via $\mathbf{\Sigma^{*}}$).
Priors on the causal effect matrix $\mathbf{A}$.
We place a spike-and-slab prior on each off-diagonal entry $a_{jh}$ of the matrix $\mathbf{A}$. The presence of a directed edge from $Y_{h}$ to $Y_{j}$ is governed by a binary inclusion variable $\gamma_{jh}$, and the effect size $a_{jh}$ is conditionally modeled as:
$a_{jh}$ $\sim\gamma_{jh}\cdot\text{N}(0,\tau_{jh})+(1-\gamma_{jh})\cdot\text{N}(0,\nu_{1}\cdot\tau_{jh}),$
$\gamma_{jh}$ $\sim\text{Bernoulli}(\rho_{jh})$
$\rho_{jh}$ $\sim\text{Beta}(a_{\rho},b_{\rho})$
$\sqrt{\tau_{jh}}$ $\sim\mathcal{C}^{+}(0,1),$
where $\mathcal{C}^{+}(0,1)$ is the standard half-Cauchy distribution. Here, we fix $\nu_{1}\ll 1$ to ensure good separation between signals and noises. Following makalic2016simple , half-Cauchy distribution can be reparameterized as:
$x\sim\mathcal{C}^{+}(0,1)\implies x^{2}\mid a\sim\text{IG}(1/2,1/a),\quad a\sim\text{IG}(1/2,1),$
where $\text{IG}(\cdot,\cdot)$ denotes the inverse-gamma distribution. This allows a closed-form Gibbs update.
Priors on the instrument effect matrix $\mathbf{B}$.
When there may be horizontal pleiotropy (the violation of the exclusion restriction assumption) or weak instruments (the violation of the relevance assumption), we do not impose fixed structural zeros in $\mathbf{B}$ but instead adopt a spike-and-slab prior for each entry of $\mathbf{B}$:
$b_{jh}$ $\sim\phi_{jh}\cdot\text{N}(0,\eta_{jh})+(1-\phi_{jh})\cdot\text{N}(0,\nu_{2}\cdot\eta_{jh}),$
$\phi_{jh}$ $\sim\text{Bernoulli}(\psi_{jh}),$
$\psi_{jh}$ $\sim\text{Beta}(a_{\psi},b_{\psi}),$
$\sqrt{\eta_{jh}}$ $\sim\mathcal{C}^{+}(0,1),$
where $\nu_{2}\ll 1$ and $\phi_{jh}$ indicates whether $X_{h}$ is a (significant) instrument for $Y_{j}$. This prior formulation encourages sparsity in $\mathbf{B}$, facilitating automatic selection of valid and relevant instruments. Under the InSIDE assumption bowden2015mendelian ; yuan2020testing that the instrument effects are independent of each other, we obtain unbiased causal effect estimation. We call MR.RGM with selection on instruments MR.RGM+. Otherwise, simple normal priors are placed on the non-zero entries of $\mathbf{B}$.
Prior on the covariate effect matrix $\mathbf{C}$.
We assume a conjugate matrix-normal prior on $\mathbf{C}\in\mathbb{R}^{p\times l}$:
$\mathbf{C}\sim\text{MN}_{p\times l}(\mathbf{0},\,\mathbf{\Sigma}^{*},\,\tau\mathbf{I}_{l}).$
Prior on the error covariance matrix $\mathbf{\Sigma^{*}}$.
The error covariance matrix $\mathbf{\Sigma}^{*}=\mathbf{D}\mathbf{D}^{T}+\mathbf{\Sigma}$ combines unmeasured confounding (via $\mathbf{D}$) and measurement noise (via $\mathbf{\Sigma}$). Since we do not know a priori the number $t$ of unmeasured confounders, we directly model $\mathbf{\Sigma}^{*}$ instead of modeling $\mathbf{D}$ and $\mathbf{\Sigma}$. We introduce a binary indicator matrix $\mathbf{Z}=(z_{jh})\in\{0,1\}^{p\times p}$ that encodes whether the off-diagonal entry $\sigma^{*}_{jh}$ is nonzero. The priors are specified as follows:
$\sigma^{*}_{jh}\sim z_{jh}\cdot\text{N}(0,\omega_{1}^{2})+(1-z_{jh})\cdot\text{N}(0,\omega_{2}^{2}),\quad j<h,$
$z_{jh}\sim\text{Bernoulli}(\pi),$
$\sigma^{*}_{jj}\sim\text{Exp}\left(\frac{\lambda}{2}\right),$
subject to the positive-definiteness of the resulting $\mathbf{\Sigma^{*}}$. The hyperparameters $\omega_{1},\omega_{2},\pi$ and $\lambda$ control the sparsity and scale of the inferred error covariance. Following wang2015 , we recommend setting $\omega_{2}\geq 0.01$ and $\omega_{1}/\omega_{2}\leq 1000$ to ensure numerical stability and encourage separation between signals and noises. The parameter $\pi$ lies between 0 and 1, with smaller values favoring sparser structures. For the exponential prior, $\lambda$ may be set to moderately large values such as 5 or 10.
Posterior Computation.
We use MCMC to draw posterior samples of the model parameters. Step-by-step updating scheme is detailed in the Supplementary Materials. Our implementation leverages Rcpp and linear algebra for efficient matrix computations.
3 Simulation Studies
In this section, we evaluate the performance of the proposed MR.RGM, using simulations designed to reflect complex causal structures commonly found in biological systems. We compare MR.RGM with several baseline methods such as MR packages OneSampleMR, mrbayes, and MendelianRandomization, which includes MR approaches based on Simple Median, Weighted Median and IVW. These baselines offer a diverse representation of current MR tools; however, none explicitly model feedback loops. For clarity, we denote the Simple Median, Weighted Median, and IVW methods from the MendelianRandomization package as MR-SimpleMedian, MR-WeightedMedian, and MR-IVW, respectively. For ablation, we also consider two variants of MR.RGM, namely, MR.RGM_NoConf, which assumes no latent confounding and was implemented by sarkar2025mrrgm , and MR.RGM+, which accounts for horizontal pleiotropy. For the implementation of MR.RGM+, rather than prespecifying instrument-trait mapping, we treat every SNP as a potential instrument for every gene/trait.
We consider three distinct scenarios with topological features commonly observed in gene regulatory networks – scale-free and small-world graphs are well-established motifs in systems biology, and horizontal pleiotropy is ubiquitous challenges in MR.
  1. 1.
    Scale-free graph with feedback loops and unmeasured confounding: A scale-free causal network with directed cycles is constructed to model reciprocal regulation. Each trait is assigned three unique instruments, and no pleiotropy is introduced.
  2. 2.
    Small-world graph with feedback loops and unmeasured confounding: Similar to Case 1, but using a small-world network topology. Each trait again receives three unique instruments, and no pleiotropy is introduced.
  3. 3.
    Small-world graph with feedback loops, unmeasured confounding, and horizontal pleiotropy: In addition to the structure in Case 2, horizontal pleiotropy is introduced by assigning one shared IV to each consecutive trait pair (traits are arbitrarily ordered).
Given the graph and IV structure, we simulate data from (15) without observed covariates, where the non-zero off-diagonal entries of $\mathbf{A}$ are sampled from $\text{Uniform}[-0.1,0.1]$, reflecting small effect sizes to capture low-signal conditions common in real data, the non-zero entries of $\mathbf{B}$ are set to be 1, the entries of $\mathbf{D}$ are sampled from $\{-1,+1\}$, the instruments and the unmeasured confounders are drawn from independent standard normal distributions, and the errors are drawn from independent centered normal distributions with variance 9.
For Case III, since the competing methods do not account for horizontal pleiotropy, we randomly assign each instrument to one of the two traits if an instrument affects two traits.
We vary sample size $n\in\{500,1000,10000,30000\}$ and number of traits $p\in\{5,10\}$, yielding 8 settings per graph structure. Each simulation is repeated 20 times. Performance is evaluated based on the following criteria:
  • Graph Recovery: To evaluate each method’s ability to recover the true causal graph, we compute area under the ROC curve (AUC), true positive rate (TPR), false discovery rate (FDR), and Matthews correlation coefficient (MCC).
  • Causal Effect Estimation: We compute maximum absolute deviation (MaxAbsDev), mean absolute deviation (MAD), and mean squared deviation (MSD) between the estimated and true causal effects among all pairs of traits.
  • Confounding Structure Recovery: We normalize the true variance-covariance matrix $\mathbf{\Sigma}^{*}=\mathbf{DD}^{T}+\mathbf{\Sigma}$ to the range of $[0,1]$ and threshold it at its empirical mean to create a true confounding structure. We then compute AUC, TPR, FDR, and MCC of our estimated confounding structure.
  • Instrument-Trait Selection Accuracy. When MR.RGM+ is applied (i.e., in Case III), we also report the AUC of the instrument-trait selection.
In the main text, we focus on visualizing the AUC for graph recovery, confounding structure recovery, and instrument-trait selection, and the MAD for causal effect estimation, for network size p=10p{=}10. The corresponding plots for network size p=5p{=}5 as well as the full set of performance tables for all metrics and scenarios are provided in the Supplementary Materials.
Results for Cases I-II.
The results for Cases I and II are similar and hence are reported together. Figures  1(a) and 2(a) show that MR.RGM consistently achieves the highest AUC in graph recovery across all sample sizes in both cases. MR.RGM_NoConf underperforms MR.RGM for large $n$, reflecting the influence of unmeasured confounding. MR-SimpleMedian, MR-WeightedMedian, MR-IVW, OneSampleMR and mrbayes, improve steadily with increasing $n$ and nearly catch up MR.RGM by n=10,000n{=}10{,}000-30,00030{,}000.
Figures 1(b) and 2(b) show that MR.RGM attains the lowest MAD for the causal effect estimation for all sample sizes. MR.RGM_NoConf is competitive at $n\in\{500,1000\}$ but become less so as $n$ grows. MR-SimpleMedian, MR-WeightedMedian, MR-IVW, OneSampleMR, and mrbayes improve with sample size and eventually surpass MR.RGM_NoConf at larger $n$, while remaining short of MR.RGM.
(a) Graph recovery: boxplots of AUC by method (x-axis) and sample size (facets; $n\in\{500,1000,10000,30000\}$).
(b) Causal effect estimation: boxplots of MAD by method (x-axis) and sample size (facets; $n\in\{500,1000,10000,30000\}$).
Figure 1: Scale-free network with feedback loops and unmeasured confounding, with network size p=10p{=}10. (a) AUC for graph recovery; (b) MAD for causal effect estimation.
(a) Graph recovery: AUC by method (x-axis) and sample size (facets; $n\in\{500,1000,10000,30000\}$).
(b) Causal effect estimation: MAD by method (x-axis) and sample size (facets; $n\in\{500,1000,10000,30000\}$).
Figure 2: Small-world network with feedback loops and unmeasured confounding, with network size p=10p{=}10. (a) AUC for graph recovery; (b) MAD for causal effect estimation.
Figures 3(a) and 3(b) show that MR.RGM recovers the confounding structure increasingly well with sample size: the median AUC approaches 1.0 for $n\geq 10{,}000$. Competing methods do not infer the confounding structure.
(a) Scale-free network: AUC across sample sizes ($n\in\{500,1000,10000,30000\}$).
(b) Small-world network: AUC across sample sizes ($n\in\{500,1000,10000,30000\}$).
Figure 3: Confounding structure recovery performance using MR.RGM under feedback loops and unmeasured confounding, with network size $p=10$. (a) Scale-free; (b) Small-world.
Results for Case III.
Figure 4(a) shows that MR.RGM+ attains the highest AUC for graph recovery across all sample sizes under horizontal pleiotropy. MR.RGM remains competitive but is modestly attenuated by horizontal pleiotropy, while MR.RGM_NoConf lags further because it omits latent confounding. MR-SimpleMedian, MR-WeightedMedian, MR-IVW, OneSampleMR, and mrbayes improve with $n$ but stay below MR.RGM+.
Figure 4(b) shows that MR.RGM+ achieves the lowest MAD across all $n$ and is the clear winner. By contrast, MR.RGM and MR.RGM_NoConf exhibit higher MAD. MR-SimpleMedian, MR-WeightedMedian, MR-IVW, OneSampleMR, and mrbayes improve with sample size and surpass MR.RGM and MR.RGM_NoConf at large $n$, yet they remain well above MR.RGM+.
(a) Graph recovery: AUC by method (x-axis) and sample size (facets; $n\in\{500,1000,10000,30000\}$).
(b) Causal effect estimation: MAD by method (x-axis) and sample size (facets; $n\in\{500,1000,10000,30000\}$).
Figure 4: Small-world network with feedback loops, unmeasured confounding, and horizontal pleiotropy, with network size p=10p{=}10. (a) AUC for graph recovery; (b) MAD for causal effect estimation.
Figure 5 shows confounding structure recovery improving with sample size. Under horizontal pleiotropy, MR.RGM+ reaches nearly perfect AUC by $n\geq 10{,}000$, whereas MR.RGM improves with $n$ but does not quite attain perfect AUC.
Figure 5: Confounding structure recovery performance using MR.RGM and MR.RGM+ in a small-world network with feedback loops, unmeasured confounding, and horizontal pleiotropy, with network size p=10p{=}10. Boxplots of AUC by method (x-axis) and sample size (facets; $n\in\{500,1000,10000,30000\}$).
Moreover, MR.RGM+ attains consistently high AUC for recovering the true instrument-trait map (Figure 6) with AUC values tightly concentrated above $0.95$ across all $n$, indicating strong selection performance even at moderate sample size.
Figure 6: Instrument-trait selection performance using MR.RGM+ in a small-world network with feedback loops, unmeasured confounding, and horizontal pleiotropy, with network size p=10p{=}10. Boxplots of AUC across sample sizes ($n\in\{500,1000,10000,30000\}$).
Scalability Analysis.
We benchmark the runtime of MR.RGM against the competing methods. We fix the number of observations at the largest value used in our simulations, n=30,000n=30{,}000, and vary the number of traits $p\in\{2,5,10,20\}$. All Bayesian methods (MR.RGM, MR.RGM_NoConf, MR.RGM+, mrbayes) use 50,000 MCMC iterations with 10,000 burn-in; for MendelianRandomization we also run 50,000 iterations. Benchmarks were executed in RStudio on an Apple M2 Pro machine (10-core CPU, 3.5 GHz) with 16 GB unified memory. Each method is run 20 times, and we report the median wall-clock runtime in seconds in Figure 7. As $p$ increases, the runtime rises for all methods, but the MR.RGM family remains practical. For example, at $p=20$, the median runtime is approximately $90$s for MR.RGM, $55$s for MR.RGM_NoConf, $233$s for MR.RGM+.
Figure 7: Median runtime versus number of traits. Lines show median wall-clock runtime (seconds) over 20 runs for each method as $p$ increases ($p\in\{2,5,10,20\}$) with fixed n=30,000n=30{,}000. Experiments were conducted in RStudio on an Apple M2 Pro machine (10-core CPU, 3.5 GHz) with 16 GB unified memory.
4 Real Data Analysis
In this section, we demonstrate the effectiveness of the proposed method by applying it to two real-world genomic datasets: the skeletal muscle samples from the GTEx v7 dataset and the B-cell samples from the OneK1K dataset. For each dataset, we consider individuals who have complete information of gene expressions, SNPs, and relevant covariates (e.g., sex and age).
We apply our algorithm to infer the causal gene regulatory networks, with the associated uncertainty quantified by the posterior probabilities of edge inclusion, and assess the presence and the structure of latent confounders. Because horizontal pleiotropy may link any SNP to multiple genes in real tissues, we use the MR.RGM+ variant, allowing the model to select the relevant SNP–gene pairs. We run the proposed MCMC with 50,000 iterations, a burn-in of 10,000, and thinning every 10 iterations, yielding 4,000 posterior samples.
4.1 Application to GTEx v7 Skeletal Muscle Tissue Data
Dataset description.
The GTEx project is a comprehensive resource designed to study the relationship between genetic variation and gene expression across multiple human tissues. We focus on the skeletal muscle samples from GTEx v7, which consists of 332 individuals with both genotype and gene expression data available. Our analysis centers around the mechanistic target of rapamycin (mTOR) signaling pathway, a key regulator of cell growth and metabolism, which has been widely studied in both physiological and pathological contexts. We select 18 genes that are well-established components or regulators of the mTOR signaling cascade:
MTOR, ERK, AMPK, PI3K, PDK1, SHIP1, VHL, GSK3B, Tel2, TSC2, MLST8, Folliculin/BHD, PKCA, PHLPP1/2, INSULIN RECEPTOR, PRAS40, FKBP12, S6K.
We extract normalized expression levels of these 18 genes for the 332 individuals, resulting in a $332\times 18$ gene expression matrix. We utilize the signif_variant_gene_pairs file provided by GTEx and identify 62 SNPs that show significant association with at least one of the 18 genes. The resulting genotype (instrument) matrix has dimensions $332\times 62$. In addition to gene expression and SNPs, we control for two individual-level covariates: sex and age. The sex variable is coded as binary (male/female), while age is discretized into ordinal bins: 20–29 $\to$ 1, 30–39 $\to$ 2, …, 70–79 $\to$ 6.
Results.
Figure  8 shows the estimated causal network. For simplicity, we only display causal relationships for which the posterior inclusion probabilities (PIPs) are over 0.85 and confounding relationships for which the PIPs are over 0.5. Blue arrows denote directed causal edges (double-headed where bidirectional), and orange curved links indicate latent confounding between gene pairs. Edge color shading indicates the PIP, reflecting the estimation uncertainty.
Figure 8: GTEx skeletal muscle mTOR signaling network. For clarity, we only display causal edges with inclusion probability $\geq 0.85$ and confounding edges with inclusion probability $\geq 0.50$. Blue arrows: causal edges (double-headed if bidirectional). Oranger curved edges: latent confounding links. Edge color shading reflects posterior support.
In Tables 1 and 2, we highlight a subset of biologically plausible causal relationships (gene regulations) and confounding structure, respectively. Generally, they align well with known molecular interactions within the mTOR signaling pathway and related regulatory cascades in muscle tissue. These results are obtained without imposing any prior knowledge about the underlying network topology, demonstrating the power of the proposed method in discovering interpretable gene regulatory relationships from genomic data.
Gene Regulation PIP Biological Interpretation
MTOR $\rightarrow$ S6K 0.871 One of the most prominent and canonical interactions in the mTORC1 pathway is the phosphorylation of ribosomal protein S6K by mTOR. This activation is essential for promoting protein synthesis and cell growth ma2009mTOR .
MTOR $\rightarrow$ PDK1 0.950 MTOR has been shown to regulate PDK1 activity, particularly under nutrient and growth factor stimulation. This regulatory axis is crucial for coordinating upstream Akt signaling and mTORC1 activation manning2007akt .
[…]
VHL $\rightarrow$ INSULIN RECEPTOR 0.984 The tumor suppressor VHL regulates hypoxia-inducible factors (HIFs), which in turn affect insulin sensitivity and receptor expression, linking VHL to metabolic regulation Gunton2020HIF .
PHLPP1/2 $\rightarrow$ PKCA 0.987 PHLPP phosphatases dephosphorylate PKC isoforms, including PKCA, playing a role in signal termination downstream of PI3K/Akt Newton2010PKC .
Tel2 $\rightarrow$ TSC2 0.878 Tel2, part of the TTT complex, stabilizes PIKK proteins such as mTOR. By influencing mTOR stability, Tel2 indirectly affects TSC2 through mTOR-mediated regulation takai2007tel2 .
MLST8 $\rightarrow$ PRAS40 0.899 MLST8 is a core component of mTORC1 and mTORC2. Through its scaffolding function, it modulates assembly of complexes that regulate PRAS40, a known mTORC1 inhibitor VanderHaar2007PRAS40 .
[…]
Table 1: A few key gene regulatory relationships identified from the GTEx v7 skeletal muscle dataset, along with their posterior inclusion probabilities (PIPs) and biological validation.
Confounding PIP Biological Interpretation
MTORTSC2 0.525 MTOR and TSC2 are part of the same regulatory axis wherein TSC2 negatively regulates mTORC1 activity. Their interaction is modulated by AMPK and insulin signaling, making their co-expression sensitive to metabolic state and upstream energy-sensing signals, a plausible source of shared confounding saxton2017mtor .
PDK1S6K 0.501 In human skeletal muscle, both PDK1 and S6K are activated downstream of insulin/PI3KAktmTOR signaling. While PDK1 can directly phosphorylate S6K1, their correlation at the systems level is more likely driven by confounding through shared upstream inputs, particularly insulin- and growth factor–mediated PI3K activity. Thus, rather than reflecting a direct causal dependency, their association in muscle tissue may arise from coordinated regulation of the anabolic signaling network alessi1998pdk1 .
Folliculin/BHDPRAS40 0.503 Both Folliculin (BHD) and PRAS40 negatively regulate mTORC1 in response to metabolic cues and are phosphorylated in response to AMPK/Akt signaling. Their activities intersect through nutrient-sensing regulatory feedbacks, suggesting coordinated regulation and shared latent influences tsun2013folliculin .
MTORFKBP12 0.501 FKBP12 is a well-characterized binding partner of MTOR, crucial for rapamycin-mediated inhibition of mTORC1. Their expression is often correlated under rapamycin treatment and nutrient signaling, indicative of potential shared upstream regulatory programs sabatini1994raft1 .
Table 2: A few biologically plausible confounding structures inferred from the GTEx v7 skeletal muscle dataset, presented with their posterior inclusion probabilities (PIPs) and supporting biological validation.
4.2 Application to OneK1K B Cell Data
Dataset description.
The OneK1K cohort is a deeply phenotyped dataset combining genotype and transcriptomic data from a large number of individuals to study immune regulatory mechanisms. In this study, we focus on the B cells from 891 individuals. B cells are essential to adaptive immunity, and the B cell receptor (BCR) signaling pathway governs key processes such as antigen recognition, proliferation, and survival.
We analyze expression data for 66 genes central to the BCR signaling cascade, including membrane receptors, kinases, adaptor proteins, transcription factors, and regulators of apoptosis. The curated genes were selected based on their involvement in distinct signaling modules:
  • Membrane receptors and proximal signaling: CD19, BCR, FGR2B, SHIP, LYN, SYK, CD22, CD45, CBP/PAG, CSK, PIR-B
  • Adaptor proteins and scaffolds: BCAP, BLNK, GRB2, LAB, BAM32, DOK1, CBL
  • PI3K-AKT-mTOR axis: P85, PI3K, PIP3, AKT, P70S6K, GSK3
[…]
  • RAS/MAPK signaling: SOS, RASGRP, RASGAP, RAS, RAP, RIAM, MEK, MEK1/2, ERK1/2, C-RAF, MEKK, JNK, P38
  • Cytoskeletal rearrangement and trafficking: EZRIN, CLATHRIN, VAV, RAC, HS1, PYK2
[…]
  • Transcriptional regulators and apoptosis mediators: CD40, ETS1, BFL1, BCL-XL, BCL6, EGR1, JUN, ATF2, CREB, MEF2C, RAPL
For these genes, we extract normalized cell-type-level gene expressions across all 891 individuals, resulting in a $891\times 66$ gene expression matrix. SNP-gene marginal association scores are obtained from the OneK1K study. We retain 847 SNPs significantly associated with at least one of the 66 genes, producing an instrument matrix of dimension $891\times 847$. We also include as covariates sex (male/female) and age (discretized: <<30 as 1, 30–39 as 2, 40–49 as 3, …, 70–79 as 6, and 80+ as 7).
Results.
To visualize the estimated causal network without clutter, we focus on the following 29 genes, whose causal relationships will be discussed in detail:
PYK2, SYK, CBL, DOK1, PIP3, AKT, CD19, PI3K, ERK1/2, JUN, NFKB, IKB, PKC, CAMK, PLCY2, VAV, RIAM, RAP, JNK, CREB, CD40, MEK1/2, MEKK, EZRIN, HS1, SHIP, FGR2B, MEF2C, RAC.
For readability, we display their regulatory relationships as nine overlapping modules (a gene may appear in more than one panel when it bridges modules) in Figure 9. Blue arrows denote causal edges; orange curved lines denote confounding links. Within each edge type, lighter lines indicate lower posterior probabilities.
In Tables 3 and 4, we highlight a subset of biologically plausible causal relationships (gene regulations) and confounding structure, respectively. Many of them align with well-established regulatory mechanisms in B cell development, signal transduction, and immune response modulation. They capture not only unidirectional regulatory relationships but also feedback loops that reflect the dynamic nature of BCR signaling.
(a) Module 1
(b) Module 2
(c) Module 3
(d) Module 4
(e) Module 5
(f) Module 6
(g) Module 7
(h) Module 8
(i) Module 9
Figure 9: OneK1K B-cell subnetworks shown in nine panels. Blue arrows denote causal edges (double-headed if bidirectional); orange curved links denote confounding between gene pairs. Within each edge type, lighter color indicates lower posterior probability. For clarity, we only display causal edges with inclusion probability $\geq 0.50$, and confounding edges with inclusion probability $\geq 0.48$. A gene may appear in multiple panels if it bridges modules.
Regulation PIP Biological Interpretation
PYK2 $\rightarrow$ SYK 0.981 PYK2 (also known as PTK2B) is a calcium-sensitive non-receptor tyrosine kinase that plays a regulatory role in immune cell signaling. In B cells, PYK2 is activated downstream of integrin and BCR engagement and has been shown to facilitate the recruitment and activation of SYK by phosphorylating ITAMs and forming signaling complexes with SYK and other adaptors. This upstream role positions PYK2 as a modulator of SYK-mediated signaling cascades, particularly in integrin-enhanced or adhesion-dependent B cell responses Tse2009 .
CBL $\rightarrow$ DOK1 0.994 CBL, an E3 ubiquitin ligase, is recruited to signaling complexes downstream of the BCR where it associates with tyrosine-phosphorylated DOK1. This interaction promotes ubiquitination and degradation of DOK1, modulating its role as a negative regulator of Ras and PI3K signaling. Through this regulation, CBL fine-tunes BCR signaling by limiting DOK1’s inhibitory effects on MAPK and survival pathways, ensuring balance between activation and attenuation of B cell responses Lupher1999Cbl .
PIP3 $\rightarrow$ AKT 0.883 In B cells, PIP3 generated by PI3K recruits AKT to the membrane via its PH domain, where it becomes activated by phosphorylation. This is central to cell survival and metabolic regulation So2012 .
CD19 $\rightarrow$ PI3K 0.949 CD19 serves as a coreceptor for BCR signaling and amplifies signal transduction by recruiting PI3K. It directly binds the p85 regulatory subunit of PI3K upon phosphorylation, enhancing PIP3 production and facilitating activation of downstream effectors such as AKT and BTK in B cells Inabe2002 .
ERK1/2 $\rightarrow$ JUN 0.927 ERK1/2, part of the MAPK cascade, phosphorylates c-JUN, a component of the AP-1 transcription factor complex. In activated B cells, this phosphorylation increases JUN transcriptional activity, promoting expression of genes involved in proliferation, differentiation, and survival Roskoski2012 .
PI3K $\rightarrow$ AKT 0.857 The PI3KAKT signaling axis is central in B cell biology. Upon activation by coreceptors or cytokines, PI3K catalyzes production of PIP3, which recruits and activates AKT. This supports survival, proliferation, and metabolic adaptation Okkenhaug2003 .
[…]
Table 3: A few key gene regulatory relationships identified from the OneK1K B-cell dataset, along with their posterior inclusion probabilities (PIPs) and biological validation.
Confounding PIP Biological Interpretation
MEK1/2ERK1/2 0.498 In B cells, both MEK1/2 (MAP2K1/2) and ERK1/2 are co-activated by RasRaf signaling following BCR stimulation. Their correlation likely reflects confounding via shared RasRaf inputs that simultaneously activate MEK1/2 and ERK1/2 kolch2000 .
MEKKJNK 0.494 MEKK (MAP3K1) and JNK (MAPK8/9) are key kinases in the MAPK signaling cascade in B cells. Their activation is strongly coordinated through shared upstream regulators, notably the CBM (CARMA1–BCL10–MALT1) complex and PKCB-mediated BCR signaling. The correlation between MEKK and JNK likely reflects confounding by common upstream effectors, as both respond to antigen receptor stimulation and stress signals Davis2000 .
EZRINHS1 0.508 EZRIN links membrane proteins to the actin cytoskeleton, while HS1 regulates actin remodeling in hematopoietic cells. Both proteins are co-regulated during immune synapse formation in B cells through calcium and PI3K signaling. Their correlation likely arises from confounding by cytoskeletal remodeling pathways Harwood2011CytoskeletonBcell .
PLCY2VAV 0.491 PLCY2 hydrolyzes PIP2 to trigger calcium release, while VAV is a GEF that activates Rac and cytoskeletal rearrangements. Both are simultaneously recruited to the BCR signalosome via SYK and BLNK, suggesting that their association reflects confounding through this shared scaffold Weber2008 .
SHIPFGR2B 0.492 SHIP is an inositol phosphatase recruited by the inhibitory receptor FGR2B. Their correlation is expected since FGR2B engagement recruits SHIP via ITIM phosphorylation by LYN. Thus, their co-regulation is driven by shared LYN kinase activity, creating a confounding structure Barlev2022 .
CREBMEF2C 0.487 CREB is activated downstream of PKA/ERK, while MEF2C responds to CAMK and MAPK signals. Both transcription factors are co-activated by BCR-induced calcium influx and MAPK cascades, producing correlated activity through shared upstream inputs rather than direct interaction Khiem2008 .
RACPYK2 0.486 RAC is a small GTPase that controls actin dynamics, while PYK2 is a focal adhesion kinase activated by calcium and integrin signaling. In B cells, both are regulated downstream of VAV and calcium signals during immune synapse formation. Their association likely reflects confounding by BCR-driven adhesion and cytoskeletal pathways Tse2009 .
Table 4: A few biologically plausible confounding structures inferred from the OneK1K B-cell dataset, presented with their posterior inclusion probabilities (PIPs) and supporting biological validation.
5 Conclusion
In this work, We introduce MR.RGM, a Bayesian multivariable, bidirectional Mendelian randomization framework that performs network-wide causal inference. The method explicitly accommodates unmeasured confounding and feedback (cyclic) structure, and it jointly selects and estimates (i) the directed causal network among traits, (ii) instrument-trait effects, and (iii) a residual covariance whose off-diagonal elements indicate latent confounding. Sparsity-inducing priors yield interpretable graphs and effect maps, while the fully Bayesian formulation provides uncertainty quantification for edges, effects, and confounding links. The inference can be carried out from sufficient statistics, so the method works seamlessly with summary-level data when individual-level records are unavailable.
Extensive simulations with realistic network structure show that MR.RGM consistently outperforms competing approaches, delivering higher graph-recovery AUC and lower effect-estimation error as sample size grows. In horizontal pleiotropic regimes, MR.RGM+ achieves the best performance for both graph and effect recovery and attains near-perfect accuracy in SNP-gene selection at moderate to large sample size. The confounding structure recovered from the estimated residual covariance attains AUCs close to one in large samples, demonstrating that the same posterior output simultaneously supports causal network estimation and latent-confounder mapping. Despite using MCMC, the framework is computationally competitive, scaling to problem sizes typical in gene regulatory network analyses (i.e., sizes of typical signaling pathways) within practical runtimes.
Applications to GTEx v7 skeletal muscle tissue and the OneK1K B-cell group further validate the approach: we recover high-probability causal edges concordant with known biology and reveal plausible latent confounding modules, all with principled uncertainty measures. Taken together, these results position MR.RGM as a powerful computational tool for multivariable MR in modern transcriptomic studies.
Supplementary Material
R package availability: The full implementation of our Bayesian MR framework is publicly available as the MR.RGM package, which includes functions for posterior inference and graph visualization. The package can be accessed via: